GNU Makeinfo

Brian J. Fox and Robert J. Chassell

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993, 1994, 1995 Free Software Foundation, Inc.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

1 What is makeinfo?

This file documents the use of the makeinfo program, versions 1.61 and later. It is an extract from the TeXinfo manual.

makeinfo is a program for converting TeXinfo files into Info files. TeXinfo is a documentation system that uses a single source file to produce both on-line information and printed output.

You can read the on-line information using Info; type info to learn about Info. See the TeXinfo manual, to learn about the TeXinfo documentation system.

1.1 Controlling Paragraph Formats

Without any special options, makeinfo fills the paragraphs that it outputs to an Info file. Filling is the process of breaking and connecting lines so that lines are the same length as or shorter than the number specified as the fill column. Lines are broken between words. With makeinfo, you can control:

• The width of each paragraph (the fill-column).

• The amount of indentation that the first line of each paragraph receives (the paragraphindentation).

1.2 Command Line Options

The following command line options are available for makeinfo.

-D var Cause var to be defined. This is equivalent to @set var in the Texinfo file.

--error-limit limit

Set the maximum number of errors that makeinfo will report before exiting (on the assumption that continuing would be useless). The default number of errors that can be reported before makeinfo gives up is 100.

--fill-column width

Specify the maximum number of columns in a line; this is the right-hand edge of a line. Paragraphs that are filled will be filled to this width. The default value for

          fill-column is 72.

--footnote-style style

Set the footnote style to style, either ‘end’ for the end node style or ‘separate’ for the separate node style. The value set by this option overrides the value set in a Texinfo file by an @footnotestyle command. When the footnote style is ‘separate’, makeinfo makes a new node containing the footnotes found in the current node. When the footnote style is ‘end’, makeinfo places the footnote references at the end of the current node.

-I dir Add dir to the directory search list for finding files that are included using the @include

command. By default, makeinfo searches only the current directory.

--no-headers

Do not include menus or node lines in the output. This results in an ascii file that you cannot read in Info since it does not contain the requisite nodes or menus; but you can print such a file in a single, typewriter-like font and produce acceptable output.

--no-split

Suppress the splitting stage of makeinfo. Normally, large output files (where the size is greater than 70k bytes) are split into smaller subfiles, each one approximately 50k bytes. If you specify ‘--no-split’, makeinfo will not split up the output file.

--no-pointer-validate
--no-validate

Suppress the pointer-validation phase of makeinfo. Normally, after a Texinfo file is processed, some consistency checks are made to ensure that cross references can be resolved, etc. See Section 1.3 [Pointer Validation], page 3.

--no-warn

Suppress the output of warning messages. This does not suppress the output of error messages, only warnings. You might want this if the file you are creating has examples of Texinfo cross references within it, and the nodes that are referenced do not actually exist.

--no-number-footnotes

Supress automatic footnote numbering. By default, makeinfo numbers each footnote sequentially in a single node, resetting the current footnote number to 1 at the start of each node.

--output file

-o file Specify that the output should be directed to file and not to the file name specified in the @setfilename command found in the Texinfo source. file can be the special token ‘-’, which specifies standard output.

--paragraph-indent indent

Set the paragraph indentation style to indent. The value set by this option overrides the value set in a Texinfo file by an @paragraphindent command. The value of indent is interpreted as follows:

• If the value of indent is ‘asis’, do not change the existing indentation at the starts of paragraphs.

• If the value of indent is zero, delete any existing indentation.

• If the value of indent is greater than zero, indent each paragraph by that number of spaces.

--reference-limit limit

Set the value of the number of references to a node that makeinfo will make without reporting a warning. If a node has more than this number of references in it, makeinfo will make the references but also report a warning.

-U var Cause var to be undefined. This is equivalent to @clear var in the Texinfo file.

--verbose

Cause makeinfo to display messages saying what it is doing. Normally, makeinfo only outputs messages if there are errors or warnings.

--version

Report the version number of this copy of makeinfo.

--help Show a summary of the commend line arguments to makeinfo.

1.3 Pointer Validation

If you do not suppress pointer-validation (by using the ‘--no-pointer-validation’ option), makeinfo will check the validity of the final Info file. Mostly, this means ensuring that nodes you have referenced really exist. Here is a complete list of what is checked:

1. If a ‘Next’, ‘Previous’, or ‘Up’ node reference is a reference to a node in the current file and is not an external reference such as to ‘(dir)’, then the referenced node must exist.

2. In every node, if the ‘Previous’ node is different from the ‘Up’ node, then the ‘Previous’ node must also be pointed to by a ‘Next’ node.

3. Every node except the ‘Top’ node must have an ‘Up’ pointer.

4. The node referenced by an ‘Up’ pointer must contain a reference to the current node in some manner other than through a ‘Next’ reference. This includes menu entries and cross references.

5. If the ‘Next’ reference of a node is not the same as the ‘Next’ reference of the ‘Up’ reference, then the node referenced by the ‘Next’ pointer must have a ‘Previous’ pointer that points back to the current node. This rule allows the last node in a section to point to the first node of the next chapter.

1.4 The Macro Facility

This chapter describes the new macro facility.

A macro is a command that you define in terms of other commands. It doesn’t exist as a command until you define it as part of the input file to . Once the command exists, it behaves much as any other command. Macros are a useful way to ease the details and tedium of writing a ‘correct’ info file. The following sections explain how to write and invoke macros.

1.4.1 How to Use Macros in

Using macros in is easy. First you define the macro. After that, the macro command is available as a normal command. Here is what a definition looks like:

@macro name{arg1, … argn} commands…

     @end macro

The arguments that you specify that the macro takes are expanded with the actual parameters used when calling the macro if they are seen surrounded by backslashes. For example, here is a definition of @codeitem, a macro which can be used wherever @item can be used, but which surrounds its argument with @code{…}.

     @macro codeitem{item}
     @item @code{\item\}
     @end macro

When the macro is expanded, all of the text between the @macro and @end macro is inserted into the document at the expansion point, with the actual parameters substituted for the named parameters. So, a call to the above macro might look like:

     @codeitem{Foo}

and would execute the following code:

     @item @code{Foo}

A special case is made for macros which only take a single argument, and which are invoked without any brace characters (i.e., ‘{’…‘}’) surrounding an argument; the rest of the line is supplied as is as the sole argument to the macro. This special case allows one to redefine some standard commands without modifying the input file. Along with the non-recursive action of macro invocation, one can easily redefine the sectioning commands to also provide index entries:

@macro chapter{name}
@chapter \name\
@findex \name\
@end macro

Thus, the text:

@chapter strlen

will expand to:

@chapter strlen
@findex strlen

1.4.2 Using Macros Recursively

Normally, while a particular macro is executing, any call to that macro will be seen as a call to a builtin command. This allows one to redefine a builtin command as a macro, and then use that command within the definition of the macro itself. For example, one might wish to make sure that whereever a term was defined with @dfn{…}, the location of the definition would appear in the concept index for the manual. Here is a macro which redefines @dfn to do just that:

@macro dfn{text}
@dfn{\text\}
@cpindex \text\
@end macro

Note that we used the builtin command @dfn within our overriding macro definition.

This behaviour itself can be overridden for macro execution by writing a special macro control command in the definition of the macro. The command is considered special because it doesn’t affect the output text directly, rather, it affects the way in which the macro is defined. One such special command is @allow-recursion.

@macro silly{arg}

@allow-recursion
\arg\
@end macro

Now @silly is a macro that can be used within a call to itself:

This text @silly{@silly{some text}} is ``some text''.

1.4.3 Using Macros As Arguments

P

Pointer validation with makeinfo . . . . . . . . . . . . . . . . . . . . 3

V

Validation of pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

How to use macros as arguments to other macros.